OpenStudy (anonymous):

Question: What is the product of all integer values of x for which /x2(squared)-9/ is a prime number?

OpenStudy (anonymous):

uh...

OpenStudy (anonymous):

thats impossible to calculate i;m guesing

OpenStudy (anonymous):

I am trying to understand WHAT the question is asking

OpenStudy (anonymous):

The answer key says the answer is 64

OpenStudy (jamesj):

is what you've written the exact wording of the question?

OpenStudy (anonymous):

x could be any number and the prime number results of the equation could be infinite

OpenStudy (mr.math):

It's not empty. Take x=4 @James.

OpenStudy (anonymous):

yes thats the exact wording

OpenStudy (jamesj):

MrM, yes. \[ x^2−9 \] always has an integer factorization because \[ x^2−9=(x−3)(x+3) \] and these factors are not 1 when x > 4. So the only integer x such that x^2 - 9 is prime is when x = 4.

OpenStudy (anonymous):

No odd number exists in the set

OpenStudy (zarkon):

x=-4 also works

OpenStudy (anonymous):

Try some even numbers

OpenStudy (jamesj):

I was assuming that x^2 - 9 had to be positive, but perhaps not.

OpenStudy (anonymous):

lol I have an arguement here for x=-4

OpenStudy (mr.math):

So the product would be -16 then?

OpenStudy (anonymous):

Prime number definition is that it should have only two factors one and itself but -4 having facctors -4 and -1 hahaha what do we do now????

OpenStudy (anonymous):

does 64 work for that equation? the answer key says the answer is somehow 64

OpenStudy (mr.math):

x^2-9 is still positive at x=-4.

OpenStudy (jamesj):

yes, I'm trying to figure out how to recover 64.

OpenStudy (zarkon):

it is 64 if you allow -2 and 2

OpenStudy (jamesj):

right, and then you get negative values of x^2 - 9

OpenStudy (zarkon):

but -5 usually is not defined as a prime number

OpenStudy (zarkon):

usually they are natural numbers

OpenStudy (mr.math):

But primes by definition are positive numbers, right?

OpenStudy (anonymous):

yeah

OpenStudy (jamesj):

Yes. But 64 would seem to imply that's what the question allows, which is unusual.

OpenStudy (anonymous):

Yeah thats what i was also telling

OpenStudy (anonymous):

there must be some range for the x values

OpenStudy (jamesj):

If the set is {-4, -2, 2, 4}, then we have -4 x -2 x 2 x 4 = 64

OpenStudy (anonymous):

Range is from [-9,infnty)

OpenStudy (jamesj):

+4 and -4 make sense. +2 and -2 are unusual.

OpenStudy (anonymous):

so what should i put for the answer?

OpenStudy (zarkon):

I would use 4(-4)=-16

OpenStudy (anonymous):

ok thanks for your time

OpenStudy (jamesj):

I agree with Z.

OpenStudy (zarkon):

I just noticed something...by /x2(squared)-9/ do you mean \[|x^2-9|\]?

OpenStudy (jamesj):

oh ... in that case ...

OpenStudy (zarkon):

if that is the case then 64 is the answer

OpenStudy (anonymous):

that is indeed the case

OpenStudy (zarkon):

then -4,4,-2,2 all work (-4)4(-2)2=64

OpenStudy (jamesj):

Right. Then our concern about "negative prime" numbers goes away. The set of such x is {-4, -2, 2, 4}

OpenStudy (anonymous):

why does that matter?

OpenStudy (jamesj):

because normally prime numbers are positive integers.

OpenStudy (anonymous):

correct

OpenStudy (mr.math):

@Zarkon: You're a problems solver, I would fan you twice if I could! :D

OpenStudy (jamesj):

without the absolute value sign, then when x = 2 or x = -2, we have \[ x^2 - 9 = -5 \] which is not a positive integer and therefore can't be a prime.

OpenStudy (jamesj):

but with the absolute value sign \[ |x^2 - 9| = |-5| = 5 \] when x = 2, -2

OpenStudy (anonymous):

so how does it come out to be 64

OpenStudy (zarkon):

we should post a sign at the beginning of the site that says don't use / /or \ \ for absolute value :)

OpenStudy (anonymous):

ah yes sorry

OpenStudy (mr.math):

We would need another site if we want to avoid wrong notations!